A firm makes items A and B and the total number of items it can make in a day is 24 It takes one hour to make an item of A and only half an hour to make an item of B. The maximum time available per day is 16 hours. The profit on an item of A is Rs. 300 and on one item of B is Rs.


1
Expert's answer
2021-04-14T01:06:42-0400

Let item A = x units

Let item B = y units

"x+y\le 24"

and,

"x+\dfrac{y}{2}\le16"


Profit, "Z = 300x+160y"

From the graph, we have points


"P(0,24) = 0+ 160\times 24 = 3840"

"Q(8,16) = 300\times16 + 0 = 4800"

"R(16,0) = 300\times 8 + 16\times160 = 4960"

"S(0,0) = 0"

"Q(8,16)" will maximise the profit.